A 4-layer is biased in the forward –blocking region with an anode-to-cathode voltage of 20V under this bias condition, the anode current is 1uA. Determine the resistance of the diode in the forward-blocking region.


1
Expert's answer
2021-10-30T11:20:03-0400

the anode to cathode voltage = 24V

the Current passing through current = 1.5 uA

As per ohm's law

V=I*R

so

R= V/I

so by putting the values of V and I in the above equation

R=24/0.0000015 ohms

R=16000000 ohms

So resistance of the diode is 16000000 ohms